OpenStudy (anonymous):

Find precise value tan -7pi/12

OpenStudy (raden):

7pi/12 = 105 degrees first, use the identity tan(-x) = - tan(x) tan(-7pi/12) = tan(-105) = - tan105

OpenStudy (raden):

because 105 = 45 + 60, so to get the exact value of tan105 deg, use the identity tan(a+b) = {tan(a)+tan(b)}/{1-tan(a)tan(b)} in this case, let a=45 and b= 60 deg

OpenStudy (raden):

it should can helps u